A jet is flying from phoenix to Houston at a speed r of 600 miles per hour. The distancevd between the two cities is about 1200 miles. Which equation can be used to find the time t it takes to make the trip?
Answer by oberobic(2304)

The standard distance formula is:
d = rt
where
d = distance
r = rate (or speed)
t = time
.
r = 600
d = 1200
.
So,
d = rt
Divide both sides by r
d/r = t
t = d/r
t = 1200/600
t = 2
.
Done